OpenStudy (anonymous):

20=(1/3)*5(1+4b+1) find b

OpenStudy (anonymous):

b = 10/4 = 5/2

OpenStudy (anonymous):

B=5/2

OpenStudy (anonymous):

how long will it take $4000 to double at 4% yearly instrest